# Hestia: Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection

### v1

@@ 3.2 Next-Best-View Hierarchical Network
Training loss functions. The look-at point and viewpoint prediction networks can be jointly trained using the RL reward due to their connection. However, our previous experiments showed no clear benefit from joint training using the RL reward. To simplify the design, we detach the gradient flow between the networks and train the look-at point prediction network with supervised learning using ground-truth targets. The ground truth look-at point $L_{t}^{\text{gt}}$ is computed as the weighted average position of the ground truth uncaptured surface $L_{t}^{\text{gt}}=\frac{\sum_{v_{i}\in U}w_{v_{i}}\,p_{v_{i}}}{\sum_{v_{i}\in U}w_{v_{i}}},$ where $U$ represents the set of voxels containing ground truth uncaptured faces, and $w_{v_{i}}=\sum_{f\in F_{v_{i}}^{\text{gt}}}1$ is defined as the total number of ground truth uncaptured faces within voxel $v_{i}$ where $F_{v_{i}}^{\text{gt}}$ is the set of ground truth uncaptured faces associated with voxel $v_{i}$. Thus, the loss function for the proposal network is formulated as $\mathcal{L}_{\text{proposal}}=\|L_{t}-L_{t}^{\text{gt}}\|^{2}.$ The loss of the viewpoint prediction network is the same as the regular RL loss $\mathcal{L}_{\text{RL}}$ which depends on the RL method used, combined with an auxiliary loss $\mathcal{L}_{\text{aux}}=\|a_{t}-a^{{}^{\prime}}_{t}\|^{2}$ to encourage the predicted action $a_{t}$ to align with the collision-free action $a^{{}^{\prime}}_{t}$. Hence, the overall loss function for Hestia is $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{RL}}+0.5\cdot\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{proposal}}.$
4 Real-world Data Collection System
Figure 4: Proposed data collection system and processes. Hestia goes beyond prior works by introducing an intelligent system using a drone with an RGB camera for data capture, Lighthouse base stations and Crazyflie for localization, and a Wifi router for wireless communication.
To demonstrate our work, we propose a real-world system (see Fig. 4), where a drone equipped with an RGB camera moves to the next-best viewpoint predicted by Hestia to capture images of an object. The system uses four HTC Lighthouse base stations and a Crazyflie 2.1 for localization and transmits images to the ground control station via wireless communication. MASt3R [14] is integrated to convert RGB images into pointmaps (e.g., depth images), and three initial viewpoints are set for real-world and virtual-world synchronization [83]. Fig. 4 illustrates four key processes of the system. Specifically, the drone captures images at three initial viewpoints, where the set waypoints process navigates the drone using a heuristic trajectory planner based on prior knowledge of the environment. Then, the capture image process commands image capture, and the receive image process transmits the image to the ground station. After capturing the initial viewpoints, the nbv prediction process predicts the next-best viewpoint based on the collected data. Four processes repeat until sufficient data is collected. For more details, please refer to Sec. J.
@@ 5.1 Experimental Setup
Hestia is trained on our processed Objaverse [61, 60] split to showcase its capability and on our Houses3K [55] split denoted as Hestia-H3K for fair comparison. We use NVIDIA IsaacLab [84] to simulate 256 scenes in parallel, with each object scaled up to 8 meters and placed in a 20$\times$20$\times$20m scene. Objects are randomly translated during training and placed at the origin and the four corners for benchmarking. An RGB-D camera is ahead of the Crazyflie drone, which starts from a random collision-free position oriented toward the object center. See Sec. I and Sec. K for more details.
5.2 Results
Train Data Method OmniObject3D Objaverse Houses3K FPS
DTU NeU-NBV [53] 77 / 48 / 73 79 / 33 / 75 78 / 38 / 75 $<$1
Houses3K ScanRL [55] 80 / 32 / 75 81 / 27 / 78 81 / 32 / 76 15
GenNBV [43] 93 / 12 / 87 91 / 14 / 86 92 / 14 / 87 26
GenNBV (Rep.) [43] 93 / 11 / 87 92 / 13 / 86 94 / 12 / 89 27
Hestia-H3K 96 /  6 / 93 95 /  8 / 91 97 /  7 / 96 25
Objaverse Hestia 97 /  4 / 93 96 /  7 / 92 97 /  6 / 94 25
@@ 5.2 Results
Table 1: Overall performance on OmniObject3D, Objaverse, and Houses3K with 30 images per object. Results are reported as mean CR (%), CD (cm), and AUC (%) over five object center positions. Both Hestia and Hestia-H3K outperform prior approaches by at least 3% in CR and AUC.
Figure 5: Reconstruction on a complex scene. Hestia and Hestia-H3K capture the scene well.
Quantitative and qualitative results. We compare Hestia with generalizable next-best-view approaches [43, 55, 53], using three metrics, coverage ratio (CR) [43], Chamfer Distance (CD), and the area under the coverage curve (AUC) [43] on point cloud reconstruction task. Online learning approaches are excluded due to non-real-time speed and limited computational resources. Evaluations are conducted on the OmniObject3D [85], Objaverse [61, 60], and Houses3K [55] test splits. Please refer to Sec. I for details of the datasets. As shown in Tab. 1, Hestia and Hestia-H3K outperform the baseline [43] by at least 3% in both CR and AUC, and their CDs are nearly half of those of the baseline [43] across all three datasets, demonstrating a substantial improvement. Surprisingly, Hestia achieves slightly better CD than Hestia-H3K, even when the latter is evaluated on its in-distribution dataset (Houses3K), further validating the benefit of using a more diverse dataset. Hestia’s superior performance on other datasets also suggests that training on larger and more diverse data enhances robustness to varying object shapes. In addition, Hestia and Hestia-H3K surpass other baselines, indicating that treating voxels as cubes, using a hierarchical action space, and mitigating spurious correlations collectively improve performance. Tab. A1, Tab. A2, and Tab. A3 show that Hestia and Hestia-H3K maintain consistent performance, with less than a 1% drop in CR across different object positions. This demonstrates the effectiveness of the look-at point predictor, derived from the proposed hierarchical structure. The reconstructed point cloud using depth images collected by Hestia, shown in Fig. 5, demonstrates that Hestia and Hestia-H3K perform well in complex scenarios with multiple objects, while other methods miss occluded parts such as the leaves of the tree or the roof soffit. Fig. A1, Fig. A2, and Fig. A3 show additional qualitative results. Hestia successfully captures diverse object shapes, ranging from daily objects to building-like structures. These results highlight Hestia’s ability to handle challenging scenarios with occlusions and fine details. In addition, our ablation study in Sec. H validates the importance and effectiveness of combining the proposed components, and the visualization results in Sec. G validate that the greedy-like training scheme can mitigate spurious correlations in our task. With all these components, Hestia achieves superior performance. Please refer to Sec. B and Sec. C for more details on the qualitative and quantitative results, and to Sec. A, Sec. F, Sec. G, Sec. I, and Sec. H for more details on the importance of the proposed components.
Real-world demonstration. We demonstrate Hestia in the proposed data collection system (see Sec. 4 and Sec. J) across several scenes (see Sec. D). Hestia actively and autonomously explores the scene starting from three initial viewpoints, using a drone equipped with an RGB camera as a mobile sensor. Fig. A5 shows that Hestia captures both shifted and non-shifted scenes well, while Fig. A6 demonstrates that Hestia performs well across diverse real-world object shapes. These results show that Hestia goes beyond prior works, which typically remain limited to simulation environments (see Sec. F), and demonstrate the feasibility of developing Hestia for real-world applications.
@@ 6 Conclusion
We present Hestia, hierarchical next-best-view exploration for systematic intelligent autonomous data collection. Hestia addresses high-dimensional viewpoints by separately predicting look-at points and camera positions. Treating voxels as cubes enables more comprehensive capture, improving coverage ratios. The greedy design mitigates spurious correlations, ensuring efficient policy learning. Trained on a more diverse dataset, Hestia is robust across varied scenarios. The integration into a real-world drone-based data collection system highlights its feasibility. We hope Hestia inspires future advancements in scalable 3D reconstruction and autonomous exploration.
7 Limitations
There are some limitations to Hestia. In particular, in real-world scenarios, Hestia, like other real-world systems, is typically constrained by hardware, budget, and safety regulations. We hope that the research community works together to prevent the misuse of artificial intelligence, strengthen the development of AI safety regulations, and reduce public concerns regarding the deployment of AI technologies. Hence, the goal of this section is to comprehensively elaborate on the limitations of Hestia, particularly in real-world applications, to help prevent potential misuse.
Figure 6: Failure cases of Hestia. Hestia may occasionally fail to capture finer 3D structures, highly self-occluded parts, nearly vertical bottom-up views, and small details on coarse object surfaces.
Due to hardware limitations (e.g., the absence of an RGB-D camera), Hestia cannot fully exhibit its potential in real-world scenarios. However, since our system integrates a depth estimator to convert RGB images into depth maps, this limitation represents a tradeoff rather than a fundamental constraint. Our experiments conducted in NVIDIA IsaacLab demonstrate the full capability of Hestia, while our real-world data collection system highlights Hestia’s robustness when a depth sensor is unavailable. Another limitation of this work is that the drone’s path planner relies on heuristic rules derived from prior knowledge of the scene. Using occupancy grids for navigation in real-world scenarios presents additional challenges, as the grid is generated from RGB images rather than RGB-D data, increasing the risk of inaccuracies. A potential improvement would be to incorporate an online path planner. However, this approach would require onboard computing for real-time planning, necessitating the use of a customized drone. Hence, a potential future step to address the above limitation is to build a customized large drone with improved payload capacity, equipped with an onboard computer and an RGB-D sensor mounted on a gimbal, to further enhance performance. Furthermore, due to drone regulations, the real-world demonstration of Hestia is conducted indoors using an indoor GPS system (e.g., HTC Lighthouse base stations). Drone policies vary across countries, and obtaining outdoor flight approvals can take up to a year in our region. Additionally, outdoor trials require significant funding, such as renting a safe test site measuring approximately 100 meters by 100 meters. As a future step, we plan to test the system with a larger customized drone in an outdoor private area to further validate its performance. Another future step is to extend Hestia to a multi-agent setting for large-scale outdoor scanning (e.g., city-scale) under power-constrained scenarios.
In addition to the above limitations, we hope that Hestia will not be misused for other types of next-best-view (NBV) tasks. In this study, we found that a greedy-like training scheme can effectively mitigate spurious correlations and is well-suited to our problem definition (see Sec. G, Sec. 3, and Sec. 5). However, the next-best-view problem is a broad research topic with varying objectives. This finding may not generalize to other NBV tasks, such as next-best-view for object tracking or next-best-view for human aesthetics, where long-term planning is more critical.
Also, although our quantitative and qualitative results (see Sec. 5 and Sec. C) demonstrate a nearly comprehensive point cloud reconstruction, we hope that Hestia will not be misused as a perfect next-best-view planner for 3D reconstruction. There are still some failure cases of Hestia (see Fig. 6). Hestia may occasionally fail to capture finer 3D structures, such as the window frames of the first row house shown in Fig. 6. It may also fail to reconstruct highly self-occluded parts, such as the pillar of the second row house. In addition, Hestia sometimes struggles to capture bottom-up views that require extreme vertical viewing angles, for example, the Lego man’s right hand and the underside of the pillar. Moreover, it may struggle to reconstruct shapes with fine details over coarse surfaces, such as tiny parts of the broccoli and anise. Adopting a multi-resolution grid structure or integrating octree-based methods to enhance the voxel grid resolution could be a potential future step to mitigate these issues.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix
@@ Acknowledgments
In this appendix, we present the theoretical grounding for treating voxels as cubes in Sec. A, additional quantitative results in Sec. B, further qualitative results in Sec. C, real-world demonstration results in Sec. D, details of the reward design in Sec. E, the novelty of the proposed components in Sec. F, the impact of spurious correlations in Sec. G, the ablation study in Sec. H, dataset details and preparation in Sec. I, the real-world system setup and associated costs in Sec. J, and training and testing details in Sec. K. We hope readers understand that the main paper is limited to 8 pages, and therefore, we have included 11 additional sections in the appendix. We have made every effort to link the main paper with the appendix to improve the reading experience.
Appendix A Theoretical Grounding
@@
Our goal is to compare the expected face visibility achieved after the scenarios terminate. Hitting each of the $k$ cubes at least once can be treated as a classical coupon–collector problem [59], whose expectation is:
$k\Bigl{(}1+\frac{1}{2}+\dots+\frac{1}{k}\Bigr{)}\approx k\ln k.$ (A1)
@@
$\Bigl{(}1-\frac{1}{6k}\Bigr{)}^{n}\;\approx\;e^{-\frac{n}{6k}},\qquad 6k\gg 1.$ (A3)
Hence, through Eq. A1 and Eq. A3, we know that after Scenario 1 stops, the ratio of the expected non-visible faces is:
$e^{-\frac{k\ln k}{6k}}=k^{-\frac{1}{6}}.$ (A4)
Appendix B Benchmark Details
In this section, we present the detailed coverage ratio (CR), Chamfer Distance (CD), and area under the coverage ratio curve (AUC) from Tab. 1, broken down by each object position setting in Tab. A1A2, and A3.
OmniObject3D Test
Method Object Center
(0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
NeU-NBV [53] 88 / 24 / 83 73 / 51 / 70 75 / 75 / 70 72 / 53 / 69 76 / 38 / 71
ScanRL [55] 87 / 21 / 80 79 / 37 / 76 84 / 21 / 78 72 / 47 / 71 76 / 35 / 72
GenNBV [43] 92 / 12 / 88 92 / 14 / 86 91 / 15 / 83 94 / 10 / 89 94 /  9 / 88
GenNBV (Rep.) [43] 93 / 10 / 87 94 /  9 / 89 92 / 12 / 85 94 / 10 / 88 92 / 11 / 85
Hestia-H3K 97 /  5 / 94 97 /  5 / 92 96 /  7 / 92 96 /  5 / 92 96 /  6 / 93
Hestia 97 /  4 / 94 97 /  4 / 93 97 /  5 / 92 97 /  4 / 93 96 /  5 / 93
Table A1: CR (%) / CD (cm) / AUC (%) comparison on the OmniObject3D test set. Both Hestia and Hestia-H3K outperform other baselines and can achieve robust performance across different object positions.
@@ Appendix B Benchmark Details
Method Object Center
(0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
NeU-NBV [53] 88 / 20 / 83 76 / 41 / 74 77 / 31 / 72 77 / 41 / 74 78 / 31 / 72
ScanRL [55] 87 / 18 / 82 80 / 31 / 78 82 / 23 / 77 77 / 37 / 76 80 / 27 / 76
GenNBV [43] 94 / 11 / 89 90 / 15 / 85 89 / 17 / 82 93 / 12 / 89 91 / 13 / 86
GenNBV (Rep.) [43] 94 / 11 / 88 92 / 12 / 88 90 / 15 / 82 92 / 12 / 88 90 / 14 / 84
Hestia-H3K 96 /  8 / 93 95 /  8 / 91 95 /  9 / 90 96 /  9 / 93 95 /  8 / 90
Hestia 96 /  7 / 93 96 /  7 / 92 96 /  6 / 91 96 /  7 / 93 96 /  8 / 91
Table A2: CR (%) / CD (cm) / AUC (%) comparison on the Objaverse test set. Both Hestia and Hestia-H3K outperform other baselines and can achieve robust performance across different object positions.
@@
Method Object Center
(0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
NeU-NBV [53] 85 / 26 / 80 73 / 52 / 71 81 / 30 / 77 72 / 53 / 70 80 / 32 / 76
ScanRL [55] 86 / 21 / 79 77 / 40 / 75 88 / 19 / 82 72 / 50 / 69 81 / 28 / 76
GenNBV [43] 94 / 10 / 89 90 / 18 / 83 91 / 16 / 84 94 / 12 / 89 93 / 12 / 89
GenNBV (Rep.) [43] 94 / 12 / 88 95 / 10 / 90 94 / 13 / 88 94 / 11 / 89 91 / 14 / 88
Hestia-H3K 98 /  6 / 96 98 /  6 / 96 97 /  8 / 95 97 /  7 / 96 97 /  7 / 95
Hestia 97 /  5 / 96 97 /  5 / 93 97 /  7 / 93 97 /  5 / 94 97 /  7 / 93
@@
Table A3: CR (%) / CD (cm) / AUC (%) comparison on the Houses3K test set. Both Hestia and Hestia-H3K outperform other baselines and can achieve robust performance across different object positions.
Both Hestia and Hestia-H3K outperform other baselines across all object position settings in the OmniObject3D [85], Objaverse [60], and Houses3K [55] test splits. Moreover, Hestia and Hestia-H3K are the only methods that demonstrate robust performance across all object position settings on all three datasets, with less than a 1% coverage ratio difference across different object configurations. Interestingly, Hestia slightly outperforms Hestia-H3K in terms of Chamfer Distance when evaluated on in-distribution data (e.g., Houses3K), highlighting the benefit of using a more diverse training set.
Appendix C Qualitative Results
In this section, we present additional qualitative results for OmniObject3D [85], Objaverse [60], and Houses3K [55] in Fig. A1, Fig. A2, Fig. A3, and Fig. A4.
Figure A1: Qualitative comparison on OmniObject3D [85]. Hestia’s viewpoints reconstruct the point clouds of real-world scanned objects accurately, while other baselines exhibit less robustness across various object shapes, often missing parts in the reconstructed point clouds.
Figure A2: Qualitative comparison on Objaverse [60]. Hestia’s viewpoints successfully reconstruct diverse and complex object shapes, while other baselines exhibit less robustness, often missing self-occluded regions or parts that require bottom-up viewpoints.
Figure A3: Qualitative comparison on Houses3K [55]. Compared to the baselines, the point clouds reconstructed from the depth maps collected by Hestia capture finer details, such as roof soffits, pillars, and windows, particularly in self-occluded areas.

### v2

@@ 3.2 Next-Best-View Hierarchical Network
Training loss functions. The look-at point and viewpoint prediction networks can be jointly trained using the RL reward due to their connection. However, our previous experiments showed no clear benefit from joint training using the RL reward. To simplify the design, we detach the gradient flow between the networks and train the look-at point prediction network with supervised learning using ground-truth targets. The ground truth look-at point $L_{t}^{\text{gt}}$ is computed as the weighted average position of the ground truth uncaptured surface:
$L_{t}^{\text{gt}}=\frac{\sum_{v_{i}\in U}w_{v_{i}}\,p_{v_{i}}}{\sum_{v_{i}\in U}w_{v_{i}}}$ (10)
where $U$ represents the set of voxels containing ground truth uncaptured faces, and $w_{v_{i}}$ is defined as the total number of ground truth uncaptured faces within voxel $v_{i}$:
$w_{v_{i}}=\sum_{f\in F_{v_{i}}^{\text{gt}}}1$ (11)
where $F_{v_{i}}^{\text{gt}}$ is the set of ground truth uncaptured faces associated with voxel $v_{i}$. Thus, the loss function for the proposal network is formulated as:
$\mathcal{L}_{\text{proposal}}=\|L_{t}-L_{t}^{\text{gt}}\|^{2}$ (12)
The loss of the viewpoint prediction network is the same as the regular RL loss $\mathcal{L}_{\text{RL}}$ which depends on the RL method used, combined with an auxiliary loss:
$\mathcal{L}_{\text{aux}}=\|a_{t}-a^{{}^{\prime}}_{t}\|^{2}$ (13)
to encourage the predicted action $a_{t}$ to align with the collision-free action $a^{{}^{\prime}}_{t}$. Hence, the overall loss function for Hestia is
$\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{RL}}+0.5\cdot\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{proposal}}$ (14)
@@ 4.1 Experimental Setup
Hestia is trained on our processed Objaverse [deitke2024objaverse, Deitke_2023_CVPR] split to showcase its full capability and on our Houses3K [peralta2020next] split denoted as Hestia-H3K for fair comparison. We use NVIDIA IsaacLab [mittal2023orbit] to randomly simulate 256 scenes in parallel, with each object scaled up to 8 meters and placed in a $\times20$$\times$20m scene. Objects are placed at the origin and the four corners for benchmarking. An RGB-D camera is ahead of the Crazyflie drone, which starts from a random collision-free position oriented toward the object center. See Secs. S8 and S10 for more details.
4.2 Overall Performance
This section addresses three questions: Q1: Is Hestia’s improvement marginal? Q2: Does the method outperform prior works [jin2023neu, guedon2023macarons, peralta2020next, chen2024gennbv] with and without large-scale training [deitke2024objaverse, Deitke_2023_CVPR]? Q3: Does large-scale training further improve performance? To answer these questions, we benchmark on three datasets [deitke2024objaverse, Deitke_2023_CVPR, wu2023omniobject3d, peralta2020next] comprising 400 diverse shapes, ensuring a comprehensive and fair comparison across methods for the point cloud reconstruction task. Given the large-scale test set, we select three generalizable baselines [jin2023neu, chen2024gennbv, peralta2020next] that do not require test-time optimization, along with one online learning approach [guedon2023macarons] for benchmarking.
Tab. 1 shows that Hestia not only outperforms prior work on all three datasets, but also achieves at least 4% and 6% gains in coverage ratio (CR) and area under the coverage ratio curve (AUC), respectively, while reducing chamfer distance (CD) by 50% compared to others. Hence, this answers Q1, showing that Hestia’s improvement is not marginal. Hestia-H3K trained on a smaller, less diverse dataset (Houses3K) still outperforms prior work, demonstrating that the improvement comes not only from large-scale diverse training but also from the proposed designs, thus answering Q2. On both OmniObject3D and Objaverse, Hestia surpasses Hestia-H3K, and even on Hestia-H3K’s own in-distribution set (Houses3K), it achieves slightly better CD, indicating that training on a larger and more diverse dataset provides additional benefits, thus answering Q3.
[…]
DTU NeU-NBV [jin2023neu] 77 48 73 79 33 75 78 38 75 78 40 74 <<1
MACARONS [guedon2023macarons] 89 17 80 85 24 76 85 26 77 86 22 78 <<1
Houses3K ScanRL [peralta2020next] 80 32 75 81 27 78 81 32 76 81 30 76 15
GenNBV [chen2024gennbv] 93 12 87 91 14 86 92 14 87 92 13 87 27
GenNBV (Rep.) [chen2024gennbv] 93 11 87 92 13 86 94 12 89 93 12 87 27
Hestia-H3K (Ours) 96 6 93 95 8 91 97 7 96 96 7 92 25
Objaverse Hestia (Ours) 97 4 93 96 7 92 97 6 94 97 6 93 25
@@ 4.2 Overall Performance
Table 1: Overall performance on OmniObject3D, Objaverse, and Houses3K with 30 images per object. Results are reported as mean CR (%), CD (cm), and AUC (%) over five object center positions. Hestia and Hestia-H3K outperform prior approaches by at least 4% and 3% in CR and by 6% and 5% in AUC, respectively, while reducing CD by nearly 50%. Interestingly, Hestia achieves slightly better CD than Hestia-H3K on in-distribution data (Houses3K).
4.3 Qualitative Comparisons
Figure 3: Point cloud reconstruction on three datasets. Hestia’s reconstructions are visibly better than those of prior approaches.
This section highlights that Hestia’s improvement also extends to visualization in the point cloud reconstruction task. Fig. 3 shows that Hestia produces more comprehensive point clouds than prior work across diverse object shapes. Specifically, prior methods fail to reconstruct parts of the teddy bear and anime figurine from OmniObject3D [wu2023omniobject3d], the underside of the stair and the cactus’s hat and hands from Objaverse [deitke2024objaverse, Deitke_2023_CVPR], and self-occluded structures such as the roof soffit or window from Houses3K [peralta2020next]. In addition, Hestia performs well on the complex scenes (see Fig. S5). This improvement is largely attributed to our design, which incorporates a hierarchical structure that better identifies missing parts of objects and models voxels as cubes rather than points, thereby preserving geometric details. More qualitative results (see Secs. S3 and S11), including failure cases, are provided in the supplementary material, and all reconstruction results are included in the supplementary video.
4.4 Translation Robustness
Figure 4: Point cloud reconstruction on three datasets. Hestia’s reconstructions are visibly better than those of prior approaches.
This section demonstrates that Hestia maintains robustness when objects are placed at different positions within the scenes. Tabs. S1, S2 and S3 provide detailed results across different object placement settings. Hestia exhibits less than 1% performance fluctuation, outperforming prior methods on all three datasets. The qualitative results (see Figs. 4 and S6) also show that Hestia’s reconstruction is more robust. These results highlight the effectiveness of the hierarchical structure, which first predicts the look-at point and then determines the capture destination.
4.5 Limited Acquisitions
This section demonstrates the suitability of Hestia for efficient 3D reconstruction, where only a limited number of views can be acquired. Tab. 2 shows that Hestia outperforms prior works by at least 12% and 5% in CR under 5-image and 15-image budgets, respectively. Specifically, Hestia achieves 92% CR with only 5 acquisitions, whereas prior work [chen2024gennbv] requires 15 images to reach comparable performance. These gains stem from the close-greedy training strategy, which not only mitigates spurious correlations over time but also enables efficient capture during inference.
Method 5 images 15 images
NeU-NBV [jin2023neu] 71 76
ScanRL [peralta2020next] 69 80
MACARONS [guedon2023macarons] 63 82
GenNBV [chen2024gennbv] 75 91
GenNBV (Rep.) [chen2024gennbv] 80 91
Hestia (Ours) 92 96
Table 2: Mean CR (%) comparison across the three datasets with limited-view acquisition. Hestia outperforms prior approaches by at least 12% and 5% with a 5-image budget and a 15-image budget. Such efficiency is crucial in real-world power-constrained settings, as robots or agents may exhaust their battery within a short time.
4.6 Inference Speed
Inference speed is critical for next-best-view planning because robots with onboard cameras must capture images for 3D reconstruction before their batteries are depleted. As shown in Tab. 1, Hestia achieves 25 FPS, which is suitable for real-time deployment. Modeling voxels as cubes rather than points does not significantly reduce inference speed. The speed also demonstrates the advantage of RL-based generalizable next-best-view approaches, since using a policy model to directly predict viewpoints removes the need to sample candidate views for prediction.
4.7 Ablation Study
This section evaluates the effectiveness of Hestia’s core components through an ablation study (see Tab. 3). We investigate three key ideas: face-aware design, a close-greedy training scheme, and a hierarchical structure. For the non-hierarchical variant, the encoded grid information is fed directly into the policy model to predict 5-DoF viewpoints without feature interpolation, since interpolation requires the look-at point. Applying the close-greedy strategy or the hierarchical structure alone eases the training process, whereas face-aware observation alone may make training more difficult but still provides complementary information. Thus, the close-greedy strategy and hierarchical structure yield stronger gains when applied individually, while combining them with face-aware observation further enhances stability and capture quality. Overall, each component contributes to performance improvements, and integrating all three delivers the best results.
Face Greedy Hier. CR $\uparrow$ CD $\downarrow$ #Pa.
Hestia 88 20 6.2M
90 17 6.2M
92 13 6.2M
94 11 4.9M
95 9 6.2M
95 8 4.9M
95 9 4.9M
96 7 4.9M
Table 3: Ablations on Objaverse [deitke2024objaverse, Deitke_2023_CVPR]. Integrating the proposed ideas yields the best performance with fewer parameters.
4.8 Application
Figure 5: Real-world demonstration of non-shifted and shifted scenes. Red boxes indicate manually initialized viewpoints, while blue boxes denote viewpoints predicted by Hestia. The results demonstrate Hestia’s feasibility in real-world environments.
This section demonstrates that Hestia is feasible for real-world deployment even without a depth camera. We use a drone equipped with an RGB camera as the mobile agent for data collection and employ a depth predictor [duisterhof2024mast3r, wang2024dust3r] to convert multi-view RGB images into depth maps. The first three images are manually selected to synchronize the real-world and virtual-world settings. As shown in Fig. 5, Hestia successfully operates in real-world scenarios for both shifted and non-shifted scenes. It is worth noting that some prior works [peralta2020next, chen2024gennbv, jin2023neu, guedon2023macarons] report only simulation results, and their real-world feasibility remains unknown. Please see Secs. S4 and S9 for more details.
@@ 5 Conclusion
We present Hestia, voxel-face-aware hierarchical next-best-view acquisition for efficient 3D reconstruction. Hestia addresses the high-dimensional action space by separately predicting look-at points and camera positions. Treating voxels as cubes enables more comprehensive capture, improving coverage ratios. The close-greedy design mitigates spurious correlations, ensuring efficient policy learning. Trained on a more diverse dataset, Hestia is robust across varied object-centric scenes. Evaluations on three datasets validate that Hestia’s improvements are not marginal. Finally, the integration into a real-world drone system highlights its feasibility. As discussed in the limitations and future steps section (Sec. S11), one important step is extending to multi-agent settings to further improve efficiency.
6 Acknowledgements
We thank Yu-Lun Liu for insightful discussions on project direction and Yi-Shan Hung for assistance with paper proofreading and demo video narration.
Supplementary Materials
@@ 6 Acknowledgements
In this supplementary material, we present the theoretical grounding for treating voxels as cubes in Sec. S1, additional quantitative results in Sec. S2, further qualitative results in Sec. S3, more real-world demonstration results in Sec. S4, details of the reward design in Sec. S5, the novelty of the proposed components in Sec. S6, the impact of spurious correlations in Sec. S7, dataset details and preparation in Sec. S8, the real-world system setup and associated costs in Sec. S9, training and testing details in Sec. S10, and limitations in Sec. S11.
S1 Theoretical Grounding
@@
Our goal is to compare the expected face visibility achieved after the scenarios terminate. Hitting each of the $k$ cubes at least once can be treated as a classical coupon–collector problem [boneh1997coupon], whose expectation is:
$k\Bigl(1+\frac{1}{2}+\dots+\frac{1}{k}\Bigr)\approx k\ln k.$ (S1)
@@
N= 6kN\;=\;6k (S2)
$\Bigl(1-\frac{1}{6k}\Bigr)^{n}\;\approx\;e^{-\frac{n}{6k}},\qquad 6k\gg 1.$ (S3)
Hence, through Eqs. S1 and S3, we know that after Scenario 1 stops, the ratio of the expected non-visible faces is:
$e^{-\frac{k\ln k}{6k}}=k^{-\frac{1}{6}}.$ (S4)
S2 Benchmark Details
In this section, we present the detailed coverage ratio (CR), Chamfer Distance (CD), and area under the coverage ratio curve (AUC) from Tab. 1, broken down by each object position setting in Tabs. S1, S2 and S3. Hestia outperforms other baselines across all object position settings in the OmniObject3D [wu2023omniobject3d], Objaverse [Deitke_2023_CVPR], and Houses3K [peralta2020next] test splits. Moreover, Hestia is the only methods that demonstrate robust performance across all object position settings on all three datasets, with less than a 1% coverage ratio difference across different object configurations. For efficient 3D reconstruction, Fig. S1 shows that Hestia outperforms other methods by nearly 10% and 5% in the first five and fifteen captures, respectively. This efficiency is especially important in real-world power-constrained scenarios, where a robot or agent may quickly exhaust its battery.
OmniObject3D Test
Method (0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
@@
NeU-NBV [jin2023neu] 88 24 83 73 51 70 75 75 70 72 53 69 76 38 71
ScanRL [peralta2020next] 87 21 80 79 37 76 84 21 78 72 47 71 76 35 72
MACARONS [guedon2023macarons] 86 19 75 93 13 86 91 14 83 94 10 88 80 29 69
GenNBV [chen2024gennbv] 92 12 88 92 14 86 91 15 83 94 10 89 94 9 88
GenNBV (Rep.) [chen2024gennbv] 93 10 87 94 9 89 92 12 85 94 10 88 92 11 85
Hestia (Ours) 97 4 94 97 4 93 97 5 92 97 4 93 96 5 93
Table S1: CR (%) / CD (cm) / AUC (%) comparison on the OmniObject3D test set. Hestia outperforms other methods and is more robust across different object position settings.
@@
Method (0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
[…]
NeU-NBV [jin2023neu] 88 20 83 76 41 74 77 31 72 77 41 74 78 31 72
ScanRL [peralta2020next] 87 18 82 80 31 78 82 23 77 77 37 76 80 27 76
MACARONS [guedon2023macarons] 88 17 79 91 16 86 91 15 83 75 42 70 78 30 64
GenNBV [chen2024gennbv] 94 11 89 90 15 85 89 17 82 93 12 89 91 13 86
GenNBV (Rep.) [chen2024gennbv] 94 11 88 92 12 88 90 15 82 92 12 88 90 14 84
Hestia (Ours) 96 7 93 96 7 92 96 6 91 96 7 93 96 8 91
Table S2: CR (%) / CD (cm) / AUC (%) comparison on the Objaverse test set. Hestia outperforms other methods and is more robust across different object position settings.
@@
Method (0, 0) (4, 4) (4, -4) (-4, 4) (-4, -4)
[…]
NeU-NBV [jin2023neu] 85 26 80 73 52 71 81 30 77 72 53 70 80 32 76
ScanRL [peralta2020next] 86 21 79 77 40 75 88 19 82 72 50 69 81 28 76
MACARONS [guedon2023macarons] 87 20 78 92 16 87 93 14 85 71 51 66 81 30 68
GenNBV [chen2024gennbv] 94 10 89 90 18 83 91 16 84 94 12 89 93 12 89
GenNBV (Rep.) [chen2024gennbv] 94 12 88 95 10 90 94 13 88 94 11 89 91 14 88
Hestia (Ours) 97 5 96 97 5 93 97 7 93 97 5 94 97 7 93
@@
Table S3: CR (%) / CD (cm) / AUC (%) comparison on the Houses3K test set. Hestia outperforms other methods and is more robust across different object position settings.
Figure S1: CR curves on three datasets. Hestia outperforms prior approaches by nearly 10% and 5% in the first five captures and the first fifteen captures, respectively. The efficiency is particularly significant in real-world power-constrained scenarios, where a robot or agent may run out of battery in a short time.
S3 Qualitative Results